in a triangle ABC how to find maximum value of cosA + cosB + cosC

Grade:12

1 Answers

sonika p
6 years ago
Maximum Value is obtained when cosA = cosB =cosC
But, A+B+C = 180
thus A=B=C=60Cos 60 =1/2
Maximum value of Cos A+cosB+cosC = 3/2